# Fractional Brownian motion with mean-density interaction

### v1

@@ IV Scaling theory
The factor $t$ accounts for the fact that the space-integral over of the integrated density $P_{\mathrm{int}}(x,t)$ increases linearly with time. As a result, the scaling function $Y$ can be normalized to unity
@@ V Results: Probability densities
where $\Gamma$ is the incomplete Gamma function (for details, see the Appendix). This function has a maximum (with a cusp) at $x=0$ and a Gaussian tail (up to a power-law prefactor) for large $x$.
@@
Despite its highly non-Gaussian form, the probability density in the interaction-dominated regime fulfills the scaling form (24) with $b(t)=x_{\mathrm{rms}}(t)=\sqrt{\langle x^{2}(t)\rangle}$, as can be seen in Fig. 6(b). The small deviations from perfect scaling collapse for the shortest time in the figure can be attributed to finite-time effects. Specifically, the force terms do not completely dominate at $t=2^{18}$, and the fractional Gaussian noise produces the small tails at large $|x|$.
We note that similar bimodal probability densities are observed, for different physical reasons, in Levy walks, certain heterogeneous diffusion processes, fractional wave equations, and the end-to-end distance of semi-flexible polymers.
VI Generalizations
So far, we have considered motion in one space dimension under the influence of a force that is proportional to the gradient of the integrated density. It is interesting to ask how the system behaves in higher dimensions and for other functional forms of the density-dependent force. The scaling theory of Sec. IV is easily generalized to $d$ space dimensions and forces that behave as the $\lambda$-th power of the gradient of $P_{\mathrm{int}}(\mathbf{x},t_{n})$.
@@ VI Generalizations
Computer simulations of FBM in higher space dimensions require a significantly larger numerical effort. For this reason, a numerical test of the generalized scaling theory is relegated to future work.
VII Conclusions
@@
Employing large-scale computer simulations as well as a one-parameter scaling theory, we have found that the behavior of one-dimensional, unbounded FBM with mean-density interaction falls in one of two regimes, depending on the value of the exponent $\alpha$ characterizing the fractional Gaussian noise. For $\alpha>4/3$, the long-time behavior is governed by the fractional Gaussian noise, and the force terms become negligibly small. Consequently, in this regime, the mean-squared displacement and the probability density agree with the corresponding quantities of noninteracting FBM for sufficiently long times. For $\alpha<4/3$, in contrast, the long-time behavior of the model is dominated by the interactions, and the fractional Gaussian noise only makes subleading contributions. As a result, the mean-squared displacement grows like $t^{4/3}$ for all $\alpha\leq 4/3$ and the probability density becomes highly non-Gaussian.
@@
Our work suggests many interesting extensions that may stimulate further research. These include the questions of higher space dimensions and nonlinear forces that we have already touched upon in Sec.  VI. It is also interesting to study what happens in applications in which the particles are attracted rather than repulsed by regions of high (integrated) density. Moreover, we expect a nontrivial interplay between the mean-density interaction and reflecting walls that confine the process to some finite or semi-finite region of space. In the current model, motivated by the growth of serotonergic fibers in the brain, the particles interact with the density of the entire trajectory ensemble, accumulated since the starting time. In addition, one can ask what changes if the interaction effect decays over time. These and other questions remain tasks for the future.
@@ Acknowledgements.
This research was supported in part by an NSF-BMBF (USA-Germany) CRCNS grant (NSF #2112862 and BMBF #STAX). The simulations were performed on the Pegasus, Foundry, and Mill clusters at Missouri S&T.
Appendix A Integrated density of FBM
@@
In this Appendix, we sketch the derivation of the expression (23) for the integrated density of (non-interacting) FBM. Equations (20) and (21) imply that the integrated density simply is a sum over the (instantaneous) probability densities,
@@
The integral over $z$ yields the incomplete Gamma function, which concludes the derivation of eq. (23). It is worth emphasizing that the expression (23) fulfills the scaling form (11) with $b(t)=x_{\mathrm{rms}}(t)=\sigma t^{\alpha/2}$. This can be seen explicitly by rewriting (23) as

### v2

@@ IV Scaling theory
The factor $t$ accounts for the fact that the space-integral (normalization) of the integrated density $P_{\mathrm{int}}(x,t)$ increases linearly with time. As a result, the scaling function $Y$ can be normalized to unity
@@
The scaling theory also allows us to estimate the crossover time $\tau_{x}$ from the initial FBM behavior to the interaction-dominated long-time behavior. According to eq. (13), the forces behave as $f\sim At^{1-\alpha}$ in the FBM regime, leading to a typical displacement $x_{\mathrm{force}}\sim At^{2-\alpha}$. The crossover occurs when this force-induced displacement overcomes the FBM displacement. This implies $\sigma\tau_{x}^{\alpha/2}\sim A\tau_{x}^{2-\alpha}$ or
$\tau_{x}\sim\left(\frac{A}{\sigma}\right)^{2/(3\alpha-4)}$ (20)
This relation holds in the interaction-dominated regime, $\alpha<4/3$. For the case of $\alpha=0.7$, the exponent in the power law (20) evaluates to 1.053-1.053- 1.053, in excellent agreement with the data in the inset of Fig. 2.
@@ V Results: Probability densities
where $\Gamma$ is the upper incomplete Gamma function (for details, see Appendix B). The function (24) has a maximum (with a cusp) at $x=0$ and a Gaussian tail (up to a power-law prefactor) for large $x$.
@@
[…]
We note that similar bimodal probability densities are observed, for different physical reasons, in Lévy walks, certain heterogeneous diffusion processes, fractional wave equations, the end-to-end distance of semi-flexible polymers, as well as regular but confined FBM.
[…]
Figure 8 presents the mean-squared displacement for $\alpha=0.7$ for particle numbers $N=1,2,16$, and 128.
Figure 8: Mean-squared displacement $\langle x^{2}\rangle$ of FBM with mean-density interaction vs. time $t$ for $\alpha=0.7$, interaction strength $A=1/40$ and different particle numbers $N$. The data are averages over 2048 ensembles ($N=1$), 1024 ensembles ($N=2$), 128 ensembles ($N=16$), and 16 ensembles ($N=128$). The dashed lines are fits of the long-time behavior for $N=1$ and $N=128$ with $\langle x^{2}\rangle=c\,t^{4/3}$.
The mean-squared displacement curve for $N=16$ is essentially indistinguishable from that for $N=128$. The same holds for $N=32$, and 64 (not shown in the figure for clarity). We conclude that the results for $N\geq 16$ represent the infinite-ensemble limit $N\to\infty$. The data for $N<16$ show some deviations from the $N\to\infty$ behavior. However, all curves (including the ones for $N<16$) asymptotically follow the same power law $\langle x^{2}\rangle\sim t^{4/3}$. This indicates that the process belongs to the same universality class for all $N$ including $N=1$.
This can be further illustrated by studying the integrated density $P_{\mathrm{int}}$ for different $N$. Figure 9 presents the integrated density at time $t=2^{20}\approx 10^{6}$ for $\alpha=0.7$ and particle numbers $N=1,2,16$, and 128.
Figure 9: Integrated density $P_{\mathrm{int}}(x,t)$ at time $t=2^{20}$ for $\alpha=0.7$, interaction strength $A=1/40$ and different particle numbers $N$. The data are averages over 4096 ensembles ($N=1$), 2048 ensembles ($N=2$), 256 ensembles ($N=16$), and 32 ensembles ($N=128$).
(As can be seen in Fig. 8, the system has reached the asymptotic regime for all $N$ at time $t=10^{6}$.) The densities for $N=16$ and $N=128$ are almost indistinguishable and represent the infinite-particle-number limit. For smaller $N$, the density broadens, in agreement with the observation that the mean-squared displacement in Fig. 8 increases with decreasing $N$ at fixed $t$. We have confirmed that the density $P_{\mathrm{int}}$ fulfills the scaling form (11) for all $N$ with $b(t)\sim t^{2/3}$, providing further evidence that the system belongs to the same universality class for all $N$. However, the shape of the scaling function $Y$ is $N$-dependent. Analogous behavior is expected for all $\alpha$ in the interaction-dominated regime $\alpha<4/3$.
VII Generalizations
In the preceding sections, we have analyzed motion in one space dimension under the influence of a force that is proportional to the gradient of the integrated density. It is interesting to ask how the system behaves in higher dimensions and for other functional forms of the density-dependent force. The scaling theory of Sec. IV is easily generalized to $d$ space dimensions and forces that behave as the $\lambda$-th power of the gradient of $P_{\mathrm{int}}(\mathbf{x},t_{n})$.
@@ VII Generalizations
Let us specifically consider the cases $d=2$ and $d=3$ for linear gradient forces, $\lambda=1$. In two dimensions, eq. (28) reduces to $\delta=1/2$. This implies that $\alpha=2\delta=1$ is the marginal value of $\alpha$ that separates the noise-dominated and interaction-dominated regimes in two dimensions. We note that this agrees with the finding that $d=2$ is the critical dimension of the so-called “true” or myopic self-avoiding random walk [38] discussed in Sec.  VIII. Analogously, for $d=3$, we find $\delta=2/5$ from equation (28), implying a critical $\alpha=0.8$.
Computer simulations of FBM with mean-density interaction in higher space dimensions require a significantly larger numerical effort. For this reason, a numerical test of the generalized scaling theory is relegated to future work.
VIII Conclusions
@@
The present paper has focused on exploring the basic properties of this stochastic process in one space dimension. Employing large-scale computer simulations as well as a one-parameter scaling theory, we have found that the behavior of unbounded FBM with mean-density interaction falls into one of two regimes, depending on the value of the exponent $\alpha$ characterizing the fractional Gaussian noise. For $\alpha>4/3$, the long-time behavior is governed by the fractional Gaussian noise, and the force terms become negligibly small. Consequently, in this regime, the mean-squared displacement and the probability density agree with the corresponding quantities of noninteracting FBM for sufficiently long times. For $\alpha<4/3$, in contrast, the long-time behavior of the model is dominated by the interactions, and the fractional Gaussian noise only makes subleading contributions. As a result, the mean-squared displacement grows like $t^{4/3}$ for all $\alpha\leq 4/3$ and the probability density becomes highly non-Gaussian.
The stochastic process introduced in this paper is related to the so-called “true” or myopic self-avoiding random walk that was introduced into the physics literature by Amit, Parisi, and Peliti [38] and further studied in Refs. [45, 46, 47]. The myopic self-avoiding random walk is defined as the problem of a particle that performs (uncorrelated) random steps on a lattice but tries to avoid already visited sites. (Note that this process is different from the usual self-avoiding walk and belongs to a different universality class [38].) Pietronero [45] showed that the mean-squared displacement of the myopic self-avoiding random walk in one dimension behaves as $\langle x^{2}\rangle\sim t^{4/3}$ which agrees with our result for uncorrelated noise, $\alpha=1$. In general dimension $d$, Pietronero found $\langle x^{2}\rangle\sim t^{4/(d+2)}$ for $d<2$ and $\langle x^{2}\rangle\sim t$ for $d\geq 2$. This agrees with the results of our generalized scaling theory in Sec. VII for $\alpha=1$ and $\lambda=1$. We thus conclude that normal Brownian motion with mean-density interaction belongs to the same universality class as the myopic self-avoiding random walk. Our process constitutes a generalization of this universality class to fractional noise, and could therefore also be called “myopic self-avoiding fractional Brownian motion”  [^3] Strictly, the myopic self-avoiding random walk is related to the $N=1$ version of our process at $\alpha=1$. However, as was shown in Sec. VI, our process belongs to the same universality class for all $N$..
@@
Our work suggests many interesting applications and extensions that may stimulate further research. These include the questions of higher space dimensions and nonlinear forces that we have already touched upon in Sec.  VII. It is also interesting to study what happens in applications in which the particles are attracted rather than repulsed by regions of high (integrated) density.
One important application is the problem that motivated the present study, viz., the system of serotonergic fibers in vertebrate brains. This application not only requires a generalization to three dimensions, it also implies simulations in a complex geometry. We expect a nontrivial interplay between the mean-density interaction and the reflecting walls that confine the process in the brain shape (or some other finite or semi-finite region of space). Specifically, we expect that the interaction will cut off the unphysical divergence of the density close to the wall observed for (non-interacting) superdiffusive FBM, potentially providing an improved description of the fiber system. Due to the high dimensionality and complex shape, these simulations are expected to require a huge numerical effort and thus remain a task for future.
Our paper may also have interesting applications in ecology. Many animals leave traces (e.g., pheromones) along their paths which then affect the behavior of other animals. This is an example of stigmergy, a mechanism of indirect interaction between agents via their environment [49, 50, 51, 52]. Our stochastic process may provide a basis for modeling stigmergic phenomena. Environmental traces often decay with time, this can be accounted for via a decaying memory kernel in the interaction, as discussed in Appendix A.
@@ Acknowledgements.
This research was supported in part by an NSF-BMBF (USA-Germany) CRCNS grant (NSF #2112862 and BMBF #STAXS). The simulations were performed on the Pegasus, Foundry, and Mill clusters at Missouri S&T.
Appendix A Interaction with finite decay time
In this Appendix, we consider a generalization of FBM with mean-density interaction to cases in which the influence of the density at previous times decays with time. This can be modeled by introducing a memory kernel into the definition of the integrated density. Specifically, we replace the total integrated density $P_{\mathrm{tot}}(x,t_{n})=NP_{\mathrm{int}}(x,t_{n})$ defined in eq. (5) with
$P_{\mathrm{tot}}(x,t_{n})$ $=$ $\sum_{m=1}^{n}e^{-(t_{n}-t_{m})/\tau_{d}}\,\sum_{j=1}^{N}\delta[x-x^{(j)}_{m}]~,\quad$ (29)
where $\tau_{d}$ is the decay time. This implies that the effect of past densities on the current force decays exponentially with time. The theory developed in the main part of the paper is recovered for $\tau_{d}\to\infty$. It is clear that the decaying memory kernel weakens the forces compared to the $\tau_{d}=\infty$ case. Thus, the results for $\alpha>4/3$ will not be affected qualitatively by the memory kernel. For long times the motion will remain to be noise-dominated with $\langle x^{2}\rangle\sim t^{\alpha}$. In contrast, the long-time behavior for $\alpha<4/3$ (which is force-dominated for $\tau_{d}=\infty$) is expected to change for finite $\tau_{d}$.
To analyze this, let us first discuss how the scaling theory of Sec. IV is modified by this change. We need to distinguish times $t\ll\tau_{d}$ and $t\gg\tau_{d}$. For short times, $t\ll\tau_{d}$, the exponential decay factor in eq. (29) can be neglected. The resulting theory is thus identical to the theory of Sec.  IV. In one dimension and for $\alpha<4/3$, the mean-squared displacement is thus expected to behave as $\langle x^{2}\rangle\sim t^{4/3}$ for $t\ll\tau_{d}$ (after the initial crossover at $\tau_{x}$ to interaction-dominated behavior discussed in Secs. III and IV).
For long times, $t\gg\tau_{d}$ the scaling theory changes. Specifically, the scaling ansatz (11) needs to be replaced by
$P_{\mathrm{int}}(x,t)=\frac{\tau_{d}}{b(t)}Y\left[\frac{x}{b(t)}\right]$ (30)
because the space integral (normalization) of $P_{\mathrm{int}}(x,t)$ is proportional to $\tau_{d}$ rather than $t$. Following the steps outlined in Sec. IV yields that the time-dependence of typical force varies as $t^{-2\delta}$ with time. If these forces dominate over the noise, they produce a typical displacement $x_{\mathrm{typ}}(t)\sim t^{1-2\delta}$. Self consistency requires that $x_{\mathrm{typ}}(t)$ behaves as $b(t)$. This implies $\delta=1-2\delta$ or $\delta=1/3$.
We therefore conclude that a new critical value of $\alpha$ emerges in the long-time regime $t\gg\tau_{d}.$ For $\alpha<2/3$, the long-time behavior remains force-dominated even after $\tau_{d}$, and the mean-squared displacement increases as $\langle x^{2}\rangle\sim t^{2/3}$. For $\alpha>2/3$, the noise dominates over the forces, leading to $\langle x^{2}\rangle\sim t^{\alpha}$. The expected behaviors of the mean-squared displacement in the different regimes of one-dimensional FBM with mean-density interaction in the presence of a finite decay time are summarized in Table 1.
$\alpha$ $t<\tau_{x}$ $\tau_{x}<t<\tau_{d}$ $t>\tau_{d}$
$\alpha>4/3$ $t^{\alpha}$ $t^{\alpha}$ $t^{\alpha}$
$2/3<\alpha<4/3$ $t^{\alpha}$ $t^{4/3}$ $t^{\alpha}$
$\alpha<2/3$ $t^{\alpha}$ $t^{4/3}$ $t^{2/3}$
Table 1: Behavior of the mean-squared displacement $\langle x^{2}\rangle$ of one-dimensional FBM with mean-density interaction in the presence of a finite decay time $\tau_{d}$. $\tau_{x}$ is the initial crossover time to interaction-dominated behavior discussed in Secs.  III and IV.
To test these predictions of the scaling theory, we have performed computer simulations of one-dimensional FBM with mean-density interaction for several values of the decay time $\tau_{d}$. We have focused on the regimes $\alpha<2/3$ and $2/3<\alpha<4/3$ because no change in behavior is expected for $\alpha>4/3$ (where the motion is already noise-dominated in the absence of the decaying memory kernel).
Figure 10 presents the mean-squared displacements for $\alpha=0.4$ and 1.0 for several values of the decay time $\tau_{d}$.
Figure 10: Mean-squared displacement $\langle x^{2}\rangle$ of FBM with mean-density interaction vs. time $t$ for $\alpha=0.4$ and 1.0 with interaction strength $A=1/5$ for several values of the decay time $\tau_{d}$. The data for noninteracting FBM are shown for comparison. The data are averages over 16 ensembles of 1024 random walkers each. The solid and dashed lines are power-law fits, for details see text.
To ensure that the three time regimes predicted by the scaling theory (see Table 1) are well separated, we have chosen a comparatively strong interaction, $A=1/5$, for these simulations. This reduces the crossover time $\tau_{x}$ from the initial FBM regime to the interaction-dominated regime. The data for $\alpha=0.4$ (shown in the upper panel of the figure) clearly feature the three predicted regimes. Initially, for time $t<\tau_{x}\approx 10$, the mean-squared displacement follows the FBM relation $\langle x^{2}\rangle\sim t^{0.4}$. At $\tau_{x}$, it crosses over to the interaction-dominated behavior $\langle x^{2}\rangle\sim t^{4/3}$ established in the main part of the paper. After the second crossover at $t=\tau_{d}$, the long-time behavior reaches the second interaction-dominated regime with $\langle x^{2}\rangle\sim t^{2/3}$. This can be seen clearly in the figure where the solid line for $\tau_{d}=10^{3}$ represents a power-law fit with the exponent 2/32/32 / 3.
The data for $\alpha=1.0$ (shown in the lower panel of Fig. 10) feature similar behavior. For $t<\tau_{x}\approx 200$, the mean-squared displacement is governed by the FBM relation $\langle x^{2}\rangle\sim t^{1.0}$. Force-dominated behavior, $\langle x^{2}\rangle\sim t^{4/3}$, is observed between $\tau_{x}$ and $\tau_{d}$. For times $t>\tau_{d}$, the behavior becomes noise-dominated again and crosses back over to $\langle x^{2}\rangle\sim t^{1.0}$. This differs for the case of $\alpha=0.4$ discussed above where the long-time behavior remains interaction-dominated.
The computer simulations thus fully confirm the predictions of the scaling theory for both $\alpha=0.4$ and $\alpha=1.0$.
Appendix B Integrated density of FBM
@@
In this Appendix, we sketch the derivation of expression (24) for the integrated density of (non-interacting) FBM. Equations (21) and (22) imply that the integrated density simply is a sum over the (instantaneous) probability densities,
@@
The integral over $z$ yields the upper incomplete Gamma function, which concludes the derivation of eq. (24). It is worth emphasizing that expression (24) fulfills the scaling form (11) with $b(t)=x_{\mathrm{rms}}(t)=\sigma t^{\alpha/2}$. This can be seen explicitly by rewriting (24) as